# BANG: Billion-Scale Approximate Nearest Neighbour Search using a Single GPU

### v2

@@ IX Conclusion
We presented BANG, a novel GPU-based ANNS method that efficiently handles large billion-size datasets exceeding GPU memory capacity, particularly using a single GPU. It keeps the graph on the CPU and brings together computation on compressed data and optimized GPU parallelization to achieve high throughput on massive datasets. BANG enables GPU-CPU computation pipelining and overlapping communication with computation. Thus, it is able to optimally utilize the GPU and CPU resources, and reduce the data transfer over the PCIe interconnect between the host and the GPU. As a result, on billion-scale datasets, it significantly outperforms the current state-of-the-art GPU-based methods, achieving 40$\times$–200$\times$ higher throughput at high recalls. A rigorous evaluation against competitive ANNS implementations demonstrated that BANG has the lowest hardware and operational cost, making it an attractive choice for practical deployment.

### v4

@@ IX Conclusion
We presented BANG, a novel GPU-based ANNS method that efficiently handles large billion-size datasets exceeding GPU memory capacity, particularly using a single GPU. It keeps the graph on the CPU and brings together computation on compressed data and optimised GPU parallelisation to achieve high throughput on massive datasets. BANG enables GPU-CPU computation pipelining and overlapping communication with computation. Thus, it is able to optimally utilise the GPU and CPU resources, and reduce the data transfer over the PCIe interconnect between the host and the GPU. As a result, on billion-scale datasets, it significantly outperforms the state-of-the-art GPU-based methods, achieving 50$\times$-400$\times$ their throughputs at high recalls. Our evaluations on a single GPU show that BANG delivers high throughput with low hardware and operational costs, making it a highly attractive option for practical deployment.